OpenStudy (anonymous):

What is the point-slope form of an equation with the slope 3/5 that passes through the point (10, −2)?

OpenStudy (aroub):

y-y1=m(x-x1) y+2=3/5(x-10) y+2=3/5x-6

OpenStudy (anonymous):

so would it be y − 10 = 3/5(x + 2)

OpenStudy (aroub):

No

OpenStudy (anonymous):

wait it would be y+10=3/5-2

OpenStudy (anonymous):

*5(x-2)

OpenStudy (aroub):

why?

OpenStudy (anonymous):

im not sure im lost

OpenStudy (aroub):

I gave you the answer.

OpenStudy (aroub):

or are you getting this from the choices you have ?

OpenStudy (anonymous):

its the choices i have but it did not make sense to me cause i think the three choices are not right

OpenStudy (anonymous):

A. y − 2 = 3 ⁄ 5(x + 10) B. y − 10 = 3 ⁄ 5(x + 2) C. y + 10 = 3 ⁄ 5(x − 2) these are the choices i was given

OpenStudy (aroub):

yes they seem to be all wrong

OpenStudy (aroub):

but i'll go with (b)

OpenStudy (anonymous):

ok thanks